# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

### v1

@@ 2 Sample a manipulable obstacle $o_{i}$ from a given distribution $f_{{\mathcal{O}}_{M}}$ ;
3 Sample a random configuration $c_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
4 if $co_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$  then
5       if $c_{i}(t)\to c^{\prime}_{i}$ is collision-free  then
6             Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update $c_{i}(t^{\prime})=c_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
7             $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup co_{i}(t)\setminus co_{i}^{\prime}$ ;
8             Given ${\mathbf{c}}(t^{\prime})$, select $\mathcal{N}(t^{\prime})$ ;
9             ${\mathbf{q}}_{\text{new}}=(\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime}))$ ;
10             return ${\mathbf{q}}_{\text{new}}$, $a(c_{i}(t),c_{i}^{\prime})$
@@ 9             ${\mathbf{q}}_{\text{new}}=(\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime}))$ ;
12             return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
@@
14       return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
@@
III-A Sampling-based Planning
We propose a sampling-based planner to address Problem 1 summarized in Alg. 1 The proposed algorithm incrementally builds a tree that explores the configuration space of the movable objects and the robot’s free space components to compute a manipulation plan $\tau$ as defined in Section II. We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}$ and ${\mathcal{E}}$ is a set of nodes and edges, respectively. The set ${\mathcal{V}}$ contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[c_{1}(t),c_{2}(t),\dots,c_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in. The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg. 1]. The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and ${\mathcal{E}}=\emptyset$ [line 1, Alg. 1]. Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg. 1]. The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found). In what follows, we present a detailed description of our method.
@@ III-A Sampling-based Planning
At every iteration $m$ of Alg. 1, we perform the following two steps. First, we sample from a given discrete mass function $f_{{\mathcal{V}}}:{\mathcal{V}}\rightarrow[0,1]$ a tree node denoted by ${\mathbf{q}}_{\text{rand}}(t)=[{\mathcal{N}}_{\text{rand}}(t),{\mathbf{c}}_{\text{rand}}(t)]\in{\mathcal{V}}$ [line 1, Alg. 1]. The mass function $f_{{\mathcal{V}}}$ defines the probability of selecting a node ${\mathbf{q}}\in{\mathcal{V}}$. Second, we sample a new node ${\mathbf{q}}_{\text{new}}$ that will be added to the tree; see Alg. 2. Specifically, given ${\mathbf{q}}_{\text{rand}}$, we compute the corresponding set of manipulable obstacles denoted by $\mathcal{O}_{M,\text{rand}}(t)$, [line 2, Alg. 2]. Then, we sample from a discrete mass function $f_{{\mathcal{O}}_{M}}:{\mathcal{O}}_{M,\text{rand}}\rightarrow[0,1]$ a manipulable obstacle $o_{i}\in\mathcal{O}_{M}(t)$ [line 2, Alg. 2]. The mass function $f_{{\mathcal{O}}_{M}}$ defines the probability of selecting a manipulable obstacle $o_{i}\in{\mathcal{O}}_{M,\text{rand}}$. Given the selected manipulable obstacle $o_{i}$, we sample a new configuration $c_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg. 2]. Designing the functions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ is deferred to Section III-B.
Figure 3: Illustration of the free-space components ${\mathcal{N}}(t)$ (left column) and ${\mathcal{N}}^{1}(t)$ (right column) in two environments.
@@
Next, we check if a collision-free path exists for the robot to move $o_{i}$ from $c_{i}(t)$ to $c_{i}^{\prime}$ [line 2, Alg. 2]. Two necessary conditions must be met for $c_{i}^{\prime}$ to admit such a collision-free path; if they are not satisfied then such a path does not exist [line 2, Alg. 2]. First, $c_{i}^{\prime}$ must not intersect other obstacles, i.e., $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus co_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$. Second, $co_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first. To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$. Correspondingly, ${\mathcal{N}}^{i}(t)$ denotes the component of ${\mathcal{P}}^{i}_{\text{free}}(t)$ containing the robot’s position ${\mathbf{p}}(t)$; see Fig. 3. Thus, the second condition requires $co_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
If the above two conditions hold, then there may exist such a collision-free path for $o_{i}$. This path is defined as a sequence of obstacle configurations, i.e., $\boldsymbol{\zeta}=c_{i}(t),c_{i}(t+1),\dots,c_{i}(t+s),\dots,c_{i}(t^{\prime})$ such that (i) $c_{i}(t^{\prime})=c_{i}^{\prime}$ and (ii) $co_{i}(t+s)\in{\mathcal{N}}^{i}(t)$, for all $s\in\{1,2,\dots,t^{\prime}-t\}$. This can be checked using any existing motion planner such as RRT or A* [3]. If such a path $\boldsymbol{\zeta}$ exists, this means the obstacle $o_{i}$ can be moved to $c_{i}^{\prime}$ giving rise to ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of the obstacle $o_{i}$ [line 2, Alg. 2]. In general, ${\mathcal{N}}^{i}(t)$ may split into multiple components once $o_{i}$ is moved to $c_{i}^{\prime}$, each adjacent to $co_{i}^{\prime}$. The robot can be in any of these components, and we randomly select one to be ${\mathcal{N}}(t^{\prime})$ [line 2, Alg. 2].
@@
Given the selected free-space component ${\mathcal{N}}(t^{\prime})$ and ${\mathbf{c}}(t^{\prime})$, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg. 2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$, respectively [lines 1, Alg. 1]. Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a(c_{i}(t),c_{i}^{\prime})$ that can enable it. We formalize this by iteratively constructing a function $A$ which returns the corresponding action for each edge in ${\mathcal{E}}$ [line 1, Alg. 1]. If a collision-free path $\boldsymbol{\zeta}$ does not exist or if $c_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg. 2].
@@
Once a node ${\mathbf{q}}_{\text{new}}$ is added to the tree, we check if ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t^{\prime})\neq\emptyset$. If that holds, it means that the robot can reach its goal configuration set ${\mathcal{P}}_{g}$. The tree nodes that satisfy this property are collected in a set ${\mathcal{X}}_{g}\subseteq{\mathcal{V}}$ [line 1, Alg. 1] that is initialized as ${\mathcal{X}}_{g}=\emptyset$ [line 1, Alg. 1]. Once the algorithm terminates, we select a tree node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$ (if it is non-empty) [lines 1-1, Alg. 1]. Then, we compute the path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(H+1)$ over ${\mathcal{G}}$ connecting ${\mathbf{q}}_{g}$ to the tree root ${\mathbf{q}}(0)$, i.e., $\tau_{{\mathbf{q}}}(0)={\mathbf{q}}(0)$, $\tau_{{\mathbf{q}}}(H+1)={\mathbf{q}}_{g}$, and $\tau_{{\mathbf{q}}}(k)\in{\mathcal{V}}$, $\forall k\in\{1,\dots,H+1\}$ [line 1, Alg. 1]. Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ modeling the sequence of manipulation actions that need to be applied to solve Problem 1 [line 1, Alg. 1]. The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any user-defined criterion such as the horizon $H$ of $\tau$ or the distance that needs to be traveled to execute $\tau$; see also Rem. III.1.
Remark III.1 (Execution of $\tau$)
Each action $\tau(k)=a(c_{i},c_{i}^{\prime})$ in $\tau$ involves moving an object $o_{i}$ from its current configuration $c_{i}$ to a new configuration $c_{i}^{\prime}$. Execution of this action should be in accordance with the corresponding path $\tau_{{\mathbf{q}}}$. This means that the robot should grasp $o_{i}$, while being in ${\mathcal{N}}$ as determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and drop-it off at $c_{i}^{\prime}$ while being in ${\mathcal{N}}^{\prime}$ determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$. According to line 2 in Alg. 2, there must exist a collision-free path for the robot to execute this movement. Also, observe that, as in [22], our planner decouples the computation of the robot’s motion from the computation of obstacle movements. Specifically, it does not determine the exact grasp poses for the objects of interest but instead tracks the free-space component of the robot.
Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).

### v2

@@ 2 Sample a manipulable obstacle $o_{i}$ from a given distribution $f_{{\mathcal{O}}_{M}}$ ;
3 Sample a random configuration ${\mathbf{c}}_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
4 if ${\bf{co}}_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and ${\bf{co}}_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus{\bf{co}}_{i}(t)$ then
5  Select ${\mathbf{p}}(t)\in{\mathcal{N}}_{\text{rand}}(t)$ and $\mathcal{N}$$^{\prime}$ and compute $\zeta$;
6  if $\exists~\zeta$ then
7     Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update ${\mathbf{c}}_{i}(t^{\prime})={\mathbf{c}}_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
8     $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup{\bf{co}}_{i}(t)\setminus{\bf{co}}_{i}^{\prime}$ ;
9     ${\mathbf{q}}_{\text{new}}=(\mathcal{N}^{\prime},{\mathbf{c}}(t^{\prime}))$ ;
10     return ${\mathbf{q}}_{\text{new}}$, $a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$
@@ 9     ${\mathbf{q}}_{\text{new}}=(\mathcal{N}^{\prime},{\mathbf{c}}(t^{\prime}))$ ;
12     return ${\mathbf{q}}_{\text{new}}=\varnothing,~a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})=\varnothing$
@@
14  return ${\mathbf{q}}_{\text{new}}=\varnothing,~a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})=\varnothing$
@@
At every iteration $m$ of Alg. 1, we perform the following two steps. First, we sample from a given discrete mass function $f_{{\mathcal{V}}}:{\mathcal{V}}\rightarrow[0,1]$ a tree node denoted by ${\mathbf{q}}_{\text{rand}}(t)=[{\mathcal{N}}_{\text{rand}}(t),{\mathbf{c}}_{\text{rand}}(t)]\in{\mathcal{V}}$ [line 1, Alg. 1]. Second, we sample a new node ${\mathbf{q}}_{\text{new}}$ that will be added to the tree; see Alg. 2. Specifically, given ${\mathbf{q}}_{\text{rand}}$, we compute the corresponding set of manipulable obstacles denoted by $\mathcal{O}_{M,\text{rand}}(t)$, [line 2, Alg. 2]. Then, we sample from a discrete mass function $f_{{\mathcal{O}}_{M}}:{\mathcal{O}}_{M,\text{rand}}\rightarrow[0,1]$ a manipulable obstacle $o_{i}\in\mathcal{O}_{M}(t)$ [line 2, Alg. 2]. Given the selected manipulable obstacle $o_{i}$, we sample a new configuration ${\mathbf{c}}_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg. 2]. Designing the functions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ is deferred to Section III-B.
@@
Next, we check if a collision-free path exists for the robot to execute $a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$, i.e., to move $o_{i}$ from ${\mathbf{c}}_{i}(t)$ to ${\mathbf{c}}_{i}^{\prime}$ [line 2, Alg. 2]. Transporting $o_{i}$ first requires the robot to move from its current configuration in ${\mathcal{N}}_{\text{rand}}$ to ${\bf{co}}_{i}(t)$ in order to grasp $o_{i}$. Such a path always exists regardless of the exact robot configuration in ${\mathcal{N}}_{\text{rand}}$ because $o_{i}$ is manipulable (i.e., ${\bf{co}}_{i}\cap{\mathcal{N}}_{\text{rand}}\neq\emptyset$) and the robot is holonomic (see Section II). Therefore, in what follows, we focus only on the existence of a collision-free path that enables the robot to move $o_{i}$ from ${\mathbf{c}}_{i}(t)$ to ${\mathbf{c}}_{i}^{\prime}$, assuming the robot is already grasping $o_{i}$. Two necessary conditions must be met for ${\mathbf{c}}_{i}^{\prime}$ to admit such a path [line 2, Alg. 2]. First, ${\mathbf{c}}_{i}^{\prime}$ must not intersect other obstacles, i.e., ${\bf{co}}_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus{\bf{co}}_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$. Second, ${\bf{co}}_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}^{\text{rand}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first. To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus{\bf{co}}_{i}(t)$. Correspondingly, ${\mathcal{N}}^{i}(t)$ denotes the component of ${\mathcal{P}}^{i}_{\text{free}}(t)$ containing the robot’s position ${\mathbf{p}}(t)$; see Fig. 3. Thus, the second condition requires ${\bf{co}}_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
Figure 3: Illustration of the components ${\mathcal{N}}(t)$ (left) and ${\mathcal{N}}^{5}(t)$ (right). Uniform vs non-uniform sampling for obstacle selection: In the left figure, only obstacles 1, 3, 4, and 5 are manipulable. Under uniform sampling, each obstacle $o_{i}$ will be selected with probability $p_{i}$ defined as $p_{1}=p_{3}=p_{4}=p_{5}=0.25$ and $p_{2}=0$. Under non-uniform sampling, if the LLM-recommended obstacle $o_{\text{LLM}}$ is $o_{3}$ and $p_{\text{obs}}=0.8$, then the corresponding probabilities become $p_{3}=0.85$, $p_{1}=p_{4}=p_{5}=0.05$, and $p_{2}=0$.
If these two conditions hold, then a collision-free path to execute $a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$ (as soon as it grasps $o_{i}$) may exist. This path is defined for the robot while carrying $o_{i}$. Thus, with slight abuse of notation, we define this path as $\zeta=[{\mathbf{p}}(t),{\mathbf{c}}_{i}(t)],[{\mathbf{p}}(t+1),{\mathbf{c}}_{i}(t+1)],\dots,[{\mathbf{p}}(t+s),{\mathbf{c}}_{i}(t+s)],\dots,[{\mathbf{p}}(t^{\prime}),{\mathbf{c}}_{i}(t^{\prime})]$ where ${\mathbf{c}}_{i}(t^{\prime})={\mathbf{c}}_{i}^{\prime}$, ${\bf{co}}_{i}(t+s)\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ for all $s\in\{1,2,\dots,t^{\prime}-t\}$, and the starting configuration ${\mathbf{p}}(t)\in{\mathcal{N}}_{\text{rand}}(t)$ is within a platform-specific threshold $\epsilon$ from $o_{i}$, enabling grasping. Note that if $o_{i}$ is moved to ${\mathbf{c}}_{i}^{\prime}$, the set ${\mathcal{N}}^{i}_{\text{rand}}(t)$ may split into multiple components, each adjacent to ${\bf{co}}_{i}^{\prime}$. After the move, the robot may be located in any of these components; we randomly select one, denoted $\mathcal{N}$$^{\prime}$, and require ${\mathbf{p}}(t^{\prime})\in{\mathcal{N}}^{\prime}$. Existence of $\zeta$ can be verified using any motion planner (e.g., RRT [1]) [line 2]. Under the considered holonomic dynamics (see Section II), the exact starting location ${\mathbf{p}}(t)$ of the robot does not affect the feasibility of transporting the object.
@@
If $\zeta$ exists, then we construct ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of $o_{i}$ [line 2, Alg. 2] and we set ${\mathcal{N}}(t^{\prime})={\mathcal{N}}^{\prime}$. Then, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg. 2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ [lines 1, Alg. 1]. [^1] We note that Alg. 12 do not specify the exact time step $t^{\prime}$ at which the node ${\mathbf{q}}_{\text{new}}$ is reached. The time stamp is included only to indicate that ${\mathbf{q}}_{\text{new}}$ will be reached at some future step. The actual time step $t$ at which a tree node $\mathbf{q}$} is reached is determined by the robot’s motion path $\xi$ used to execute a plan $\tau$ that passes through $\mathbf{q}$}; computation of $\xi$ is discussed later. Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$ that can enable it; see Fig. 2. We formalize this by iteratively constructing a function $A$ which returns the corresponding action for each edge in $\mathcal{E}$} [line 1, Alg. 1]. If a collision-free path $\zeta$ does not exist or if ${\mathbf{c}}_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg. 2].
@@
Once a node ${\mathbf{q}}_{\text{new}}$ is added to the tree, we check if ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t^{\prime})\neq\emptyset$. If that holds, it means that the robot can reach its goal configuration set ${\mathcal{P}}_{g}$. The tree nodes that satisfy this property are collected in a set ${\mathcal{X}}_{g}\subseteq{\mathcal{V}}$ [line 1, Alg. 1] that is initialized as ${\mathcal{X}}_{g}=\emptyset$ [line 1, Alg. 1]. Once the algorithm terminates, we select a tree node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$ (if it is non-empty) [lines 1-1, Alg. 1]. Then, we compute the path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(H+1)$ over $\mathcal{G}$} connecting ${\mathbf{q}}_{g}$ to the tree root ${\mathbf{q}}(0)$, i.e., $\tau_{{\mathbf{q}}}(0)={\mathbf{q}}(0)$, $\tau_{{\mathbf{q}}}(H+1)={\mathbf{q}}_{g}$, and $\tau_{{\mathbf{q}}}(k)\in{\mathcal{V}}$, $\forall k\in\{1,\dots,H+1\}$ [line 1, Alg. 1]. Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ [line 1, Alg. 1]. The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any criterion (e.g., the horizon $H$ of $\tau$).
Once $\tau=\tau(0),\dots,\tau(k),\dots\tau(H)$ is computed, a robot motion path $\xi$, defined as a sequence of robot configurations $\mathbf{p}$}, can be constructed sequentially across $k=0,\dots,H$. Each action $\tau(k)=a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$ involves moving an object $o_{i}$ from its current configuration ${\mathbf{c}}_{i}$ to a new configuration ${\mathbf{c}}_{i}^{\prime}$. Execution of $\tau(k)$ follows the corresponding path $\tau_{{\mathbf{q}}}$: the robot moves towards $o_{i}$ to grasp it while in $\mathcal{N}$}, determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and then drops it at ${\mathbf{c}}_{i}^{\prime}$ while in $\mathcal{N}$$^{\prime}$, determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$. This path can be computed using any existing motion planner (e.g., RRT). A motion path executing $\tau(k)$ exists because (i) the robot is holonomic and (ii) by construction of Algorithm 2 (i.e., existence of collision-free paths $\zeta$ for each $\tau(k)$). Concatenating these motion paths for all $k$ yields the full robot motion path $\xi$ [line 1, Alg. 1].